OpenStudy (anonymous):

Solve the differential equation the product of 1 over 8 times x and dy dx equals y times the square root of the quantity 4 times x squared minus 1.

OpenStudy (anonymous):

@jim_thompson5910 @VeritasVosLiberabit

OpenStudy (anonymous):

Separated the variables.

OpenStudy (anonymous):

Then took the integral.

OpenStudy (anonymous):

Answer should be C no?

OpenStudy (anonymous):

That does seem to be the only answer that makes sense

OpenStudy (anonymous):

@jim_thompson5910 can you confirm?

jimthompson5910 (jim_thompson5910):

one moment

OpenStudy (anonymous):

No problem.

jimthompson5910 (jim_thompson5910):

\[\Large \frac{1}{8x} \frac{dy}{dx} = y\sqrt{4x^2-1}\] \[\Large \frac{dy}{y} = 8x\sqrt{4x^2-1}dx\] \[\Large \int \frac{dy}{y} = \int 8x\sqrt{4x^2-1}dx\] \[\Large \ln(|y|) = \frac{2}{3}\sqrt{(4x^2-1)^3}+C\] \[\Huge |y| = e^{{}^{\frac{2}{3}\sqrt{(4x^2-1)^3}+C}}\] \[\Huge |y| = e^{{}^{\frac{2}{3}\sqrt{(4x^2-1)^3}}}*e^{C}\] \[\Huge |y| = e^{C}*e^{{}^{\frac{2}{3}\sqrt{(4x^2-1)^3}}}\] \[\Huge |y| = Ce^{{}^{\frac{2}{3}\sqrt{(4x^2-1)^3}}}\] If y >= 0, then we can drop the absolute values \[\Huge y = Ce^{{}^{\frac{2}{3}\sqrt{(4x^2-1)^3}}}\] so the answer is actually A

OpenStudy (anonymous):

Oooohhhh you bring the C along, god bless.

OpenStudy (anonymous):

Well then, okay, thanks @jim_thompson5910

jimthompson5910 (jim_thompson5910):

yes, `C` is a constant, so `e^C` is a constant too. Which is why I was able to replace `e^C` with just `C` to make things simpler.